OpenStudy (anonymous):

can someone help me solve this?!?!? \ What is the value of x in the solution to the following system of equations? x − 2y = 2 3x + y = 6

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

multiply the second equation by 2

OpenStudy (anonymous):

6x+2y=12?

Vocaloid (vocaloid):

right, now let's move to the next step our equations are: x-2y = 2 6x+2y = 12 add the two equations together

OpenStudy (anonymous):

6x+4y=14

Vocaloid (vocaloid):

not quite, can you try adding 6x + x?

OpenStudy (anonymous):

7x+4y=14?

Vocaloid (vocaloid):

close but not quite -2y + 2y = ?

OpenStudy (anonymous):

y

Vocaloid (vocaloid):

no, let's try this a bit differently -2 + 2 = ?

OpenStudy (anonymous):

0

Vocaloid (vocaloid):

right! -2+2 = 0 so what is -2y+2y = ?

OpenStudy (anonymous):

0y

Vocaloid (vocaloid):

right! 0y, which is the same as 0 7x + 0 = 14, what is x?

OpenStudy (anonymous):

7

Vocaloid (vocaloid):

no, please try again 7*x = 14 what is the value of x?

OpenStudy (anonymous):

2?

Vocaloid (vocaloid):

great! x = 2

OpenStudy (anonymous):

so thats the answer?

Vocaloid (vocaloid):

please look at the question again...

OpenStudy (anonymous):

yes thats the answer